### 11 Jan 2024 v1

# End-to-end Learnable Clustering for Intent Learning in Recommendation

Abstract.
Mining users’ intents plays a crucial role in sequential recommendation. The recent approach, ICLRec, was introduced to extract underlying users’ intents using contrastive learning and clustering. While it has shown effectiveness, the existing method suffers from complex and cumbersome alternating optimization, leading to two main issues. Firstly, the separation of representation learning and clustering optimization within a generalized expectation maximization (EM) framework often results in sub-optimal performance. Secondly, performing clustering on the entire dataset hampers scalability for large-scale industry data. To address these challenges, we propose a novel intent learning method called ELCRec, which integrates representation learning into an End-to-end Learnable Clustering framework for Recommendation. Specifically, we encode users’ behavior sequences and initialize the cluster centers as learnable network parameters. Additionally, we design a clustering loss that guides the networks to differentiate between different cluster centers and pull similar samples towards their respective cluster centers. This allows simultaneous optimization of recommendation and clustering using mini-batch data. Moreover, we leverage the learned cluster centers as self-supervision signals for representation learning, resulting in further enhancement of recommendation performance. Extensive experiments conducted on open benchmarks and industry data validate the superiority, effectiveness, and efficiency of our proposed ELCRec method. Code is available at: https://github.com/yueliu1999/ELCRec.
Sequential Recommendation, Intent Learning, Clustering Algorithm, Contrastive Learning
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; xxxxxxxxx; xxxxxxxxxprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Personalizationccs: Information systems Clusteringccs: Information systems Social recommendationccs: Information systems Personalization
1. Introduction
Sequential recommendation (SR) aims to accurately recommend items to users by evaluating historical user interactions. In recent years, due to the powerful representation learning ability of deep neural networks (DNN), DNN-based sequential recommendation methods(SASRec, ; BERT4Rec, ; RRN, ; CosRec, ; TGSRec, ; TiSASRec, ; DSSRec, ; CL4Rec, ; S3_rec, ; ICLRec, ) have attracted people’s attention High level of focus and commitment to recommended performance achieved.
Recently, researchers (ICLRec, ) claimed that a core factor in the effectiveness of recommender systems is capturing and leveraging users’ underlying intentions for items, as shown in Figure 1. To this end, a general learning paradigm called ICLRec (ICLRec, ) is proposed to discover users’ latent intentions for SR. Specifically, $k$-means clustering is used to capture potential intentions in user behavior sequences and identify cluster centers as intentions.
Figure 1. Intent learning in recommendation.
Although the recommendation performance is promising, the complex and tedious alternating optimization in existing methods leads to two problems, as shown in Fig. 2. (1) First, in the generalized expectation maximization (EM) framework, representation learning and clustering optimization are separated, which can easily lead to suboptimal performance. Additionally, this separate and complex process increases implementation difficulty while reducing repeatability. (2) Secondly, ICLRec needs to apply a clustering algorithm to the entire data, which cannot optimize the clustering distribution on small batches of data, easily causing problems of insufficient memory or long running time. Therefore, scalability is limited, especially for large real-world data.
Figure 2. Overview and issues of ICLRec (ICLRec, ).
To solve these problems, a novel intention learning method called ELCRec is proposed, as shown in Figure 3. The basic idea of this method is to better capture user intent by unifying sequential representation learning in a End-to-end Llearnable Clustering framework, thereby achieving better underlining recommendedrecommended. Specifically, the user’s behavioral process is first embedded into the latent space. Based on the behavioral embeddings, the cluster centers are then initialized as learnable neural network parameters. Furthermore, we design a clustering loss to guide the network to replace different cluster centers and merge homogeneous samples to cluster centers. This loss function allows our method to optimize the cluster distribution on small batches of data, improves the scalability of our method and solves problem (2). Furthermore, to further improve recommendation performance, the learned cluster centers are assigned as self-monitoring signals for sequential representation learning. The optimizer minimizes both clustering loss and recommendation loss simultaneously. These settings unify representation learning and cluster optimization in an end-to-end framework, improving recommendation performance and simplifying deployment. Therefore, problem (1) is also solved.
The main contributions of this paper are summarized as follows.
  • A intent learning method termed ELCRec is proposed to better mine users’ underlying intents via unifying representation learning and clustering optimization.
  • An end-to-end learnable clustering module is presented to optimize clustering distribution via mini-batch data, thus improving the scalability and convenience of deployment.
  • Extensive experiments demonstrate the capability of ELCRec from six perspectives, including superiority, effectiveness, efficiency, sensitivity, convergence, and visualization.
2. Related Work
2.1. Sequential Recommendation
Sequential Recommendation (SR) poses a significant challenge as it strives to accurately capture users’ evolving interests and recommend relevant items by learning from their historical behavior sequences. In the early stages, classical techniques such as Markov Chains and matrix factorization have been employed to assist models (Fossil, ; FM, ; FPMC, ) in learning patterns from past transactions.
In recent years, the field of deep learning has garnered significant attention and has demonstrated promising advancements across various domains including vision and language. Inspired by the remarkable success of Deep Neural Networks (DNNs), researchers have developed a range of deep Sequential Recommendation (SR) methods. For instance, Caser (Caser, ) leverages Convolutional Neural Networks (CNNs) (CNN, ) to embed item sequences into an ”image” representation over time, enabling the learning of sequential patterns through convolutional filters. Similarly, GRU4Rec (GRU4Rec, ) utilizes Recurrent Neural Networks (RNNs) (RNN, ), specifically the Gated Recurrent Unit (GRU), to model entire user sessions. The Transformer architecture (transformer, ) has also gained significant popularity and has been extended to the recommendation domain. For example, SASRec (SASRec, ) employs a unidirectional Transformer to model users’ behavior sequences, while BERT4Rec (BERT4Rec, ) utilizes a bidirectional Transformer to encode behavior sequences from both directions. To enhance the time and memory efficiency of Transformer-based SR models, LSAN (LSAN, ) introduces aggressive compression techniques for the original embedding matrix. Addressing the cold-start issue in SR models, ASReP (ASReP, ) proposes a pre-training and fine-tuning framework. Furthermore, researchers have explored the layer-wise disentanglement of architectures (ADT, ) and introduced novel modules like the Wasserstein self-attention module in STOSA (STOSA, ) to model item-item position-wise relationships. In addition to Transformers, graph neural networks (GNN_1, ; GNN_2, ; GNN_3, ; GNN_4, ) and multilayer perceptrons (MLP_1, ; MLP_2, ; MLP_3, ) have also found applications in recommendation systems.
Recently, Self-Supervised Learning (SSL), particularly contrastive learning, has gained popularity due to its ability to learn patterns from large-scale unlabeled data. As a result, SSL-based SR models have been increasingly introduced. For instance, in CoSeRec (CoSeRec, ), Liu et al. propose two informative augmentation operators that leverage item correlations to generate high-quality views. They then utilize contrastive learning to bring positive sample pairs closer together while pushing negative pairs apart. Another contrastive SR method, ECL-SR (ECL, ), is designed to ensure that the learned embeddings are sensitive to invasive augmentations while remaining insensitive to mild augmentations. Additionally, DCRec (DCRec, ) addresses the issue of popularity bias through a debiased contrastive learning framework. Techniques such as hard negative mining (GNNO, ; gSASRec, ) have also proven beneficial for recommendation systems.
In recent studies, it has been argued by researchers that users’ intents play a pivotal role in recommendation systems. For instance, Chen et al. (ICLRec, ) introduce a general framework called ICLRec, which utilizes $k$-Means clustering to extract users’ underlying intents. However, the alternating optimization process in ICLRec can be intricate and burdensome, thus hindering scalability and reproducibility. To address this limitation, this paper presents an end-to-end learnable clustering framework for intent learning in recommendation systems.
2.2. Intent Learning for Recommendation
The preferences of users towards items are implicitly reflected in their intents. Recent studies (ComiRec, ; MIND, ; Intention-aware, ; MCPRNs, ; MITGNN, ; ICM-SR, ) have highlighted the significance of users’ intents in enhancing the performance of recommendation systems. For instance, MCPRN (MCPRNs, ) introduces a mixture-channel method to model subsets of items with multiple purposes. Inspired by capsule networks (CapsNet, ), MIND (MIND, ) utilizes dynamic routing to learn users’ multiple interests. Furthermore, ComiRec (ComiRec, ) employs a multi-interest module to capture diverse interests from user behavior sequences, while the aggregation module combines items from different interests to generate overall recommendations. MITGNN (MITGNN, ) treats intents as translated tail entities and learns embeddings using graph neural networks. Pan et al. (ICM-SR, ) propose an intent-guided neighbor detector to identify relevant neighbor sessions, followed by a gated fusion layer that adaptively combines the current session with the neighbor sessions. Ma et al. (DSSRec, ) aim to disentangle the intentions underlying users’ behaviors and construct sample pairs within the same intention. The ASLI method (ASLI, ) incorporates a temporal convolutional network layer to extract latent users’ intents. Chang et al. (chang2023latent, ) formulate users’ intents as latent variables and infer them based on user behavior signals using the Variational Auto Encoder (VAE) (VAE, ). To mitigate noise caused by data augmentations in contrastive SR models, IOCRec (IOCRec, ) proposes building high-quality views at the intent level. More recently, a general latent learning framework called ICLRec (ICLRec, ) is introduced, which utilizes contrastive learning and $k$-Means clustering (kmeans, ). It learns the intent variable from all users’ behavior sequences and maximizes mutual information among users with similar intents.
2.3. Clustering Algorithm
Clustering is a fundamental and challenging task that aims to group samples into distinct clusters without supervision. By leveraging the power of unlabeled data, clustering algorithms have found applications in various domains, including computer vision (image_clustering, ), natural language processing (text_clustering, ), graph learning (SCGC, ), and recommendation systems (ICLRec, ). In the early stages, several traditional clustering methods (kmeans, ; spectral_clustering, ; GMM, ; DBSCAN, ; DPCA, ) were proposed. For instance, the classical $k$-Means clustering (kmeans, ) iteratively updates cluster centers and assignments to group samples. Spectral clustering (spectral_clustering, ) constructs a similarity graph and utilizes eigenvalues and eigenvectors to perform clustering. Additionally, probability-based Gaussian Mixture Models (GMM) (GMM, ) assume that the data distribution is a mixture of Gaussian distributions and estimate parameters through maximum likelihood. In contrast, density-based methods (DBSCAN, ; DPCA, ; meanshift, ) overcome the need for specifying the number of clusters as a hyperparameter.
In recent years, the impressive performance of deep learning has sparked a growing interest in deep clustering (CC, ; DeepDPM, ; deep_clustering_survey_4, ; SL, ; CoKe, ; PCL, ). For instance, Xie et al. propose DEC (DEC, ), a deep learning-based approach for clustering. They initialize cluster centers using $k$-Means clustering and optimize the clustering distribution using a Kullback-Leibler divergence clustering loss (DEC, ). IDEC (IDEC, ) improves upon DEC by incorporating the reconstruction of original information from latent embeddings. JULE (JULE, ) and DeepCluster (DeepCluster, ) both adopt an iterative approach, where they update the deep network based on learned data embeddings and clustering assignments. SwAV (SwAV, ), an online method, focuses on clustering data and maintaining consistency between cluster assignments from different views of the same image. DINO (DINO, ) introduces a momentum encoder to address representation collapse. Additionally, SeCu (SeCu, ) proposes a stable cluster discrimination task and a hardness-aware clustering criterion. While deep clustering has been extensively applied to image data, it is also utilized in graph clustering (DCRN, ; DAEGC, ; ARGA, ; HSAN, ) and text clustering (text_clustering, ; text_2, ; text_3, ; text_4, ). However, the application of clustering-based recommendation (ICLRec, ) is relatively uncommon. Leveraging the unsupervised learning capabilities of clustering could bring significant benefits to intent learning in recommendation systems.
Figure 3. Overview of our proposed ELCRec. At first, the encoder embeds the users’ historical behavior sequences into the latent space. Based on the embeddings, we initialize the cluster centers as the learnable neural parameters (tensor with gradient). In addition, the clustering distribution can be optimized by the proposed clustering loss with the mini-batch data, thus improving the scalability. Besides, the cluster-assisted contrastive learning is presented to assist the sequential representation learning. By these settings, the clustering optimization and representation learning are unified into an end-to-end framework, improving the recommendation performance and convenience of deployment.
3. Methodology
In this section, we will present our proposed framework, ELCRec. Firstly, we provide the necessary notations and define the task. Next, we analyze and identify the limitations in existing intent learning methods. Finally, we propose solutions to address these challenges.
3.1. Basic Notation
In a recommendation system, $\mathcal{U}$ denotes the user set, and $\mathcal{V}$ denotes the item set. For each user $u\in\mathcal{U}$, the historical behaviors are described by a sequence of interacted items $S^{u}=[s_{1}^{u},s_{2}^{u},...,s_{t}^{u},...,s_{|S^{u}|}^{u}]$. $S^{u}$ is sorted by time. $|S^{u}|$ denotes the interacted items number of user $u$. $s^{u}_{t}$ denotes the item which is interacted with user $u$ at $t$ step. In practice, during sequence encoding, the historical behavior sequences are limited with a maximum length $T$ (GRU4Rec, ; ICLRec, ; SASRec, ; Caser, ). The sequences truncated and remain the most recent $T$ interacted items if the length is greater than $T$. Besides, the shorter sequences are filled with “padding” items on the left until the length is $T$. After sequence encoding, the encoder outputs the embedding $\mathbf{S}^{u}\in\mathbb{R}^{|S^{u}|\times d}$ for sequence $S^{u}$. The basic notations are listed in Table 6.
3.2. Task Definition
Given the user set $\mathcal{U}$ and the item set $\mathcal{V}$, the recommendation system aims to precisely model the user interactions and recommend items to users. Take user $u$ for an example, the sequence encoder firstly encodes the user’s historical behaviors $S^{u}$ to the latent embedding $\mathbf{E}^{u}$. Then, based on the historical behavior embedding, the target of recommendation task is to predict the next item that is most likely interacted by user $u$ at $|S^{u}|+1$ step.
3.3. Problem Analyses
This paper aims to enhance the recommendation performance by mining the users’ underlying intents. The details of previous intent learning methods refer to related work at Section 2.2. The most similar work to our method is ICLRec (ICLRec, ). In this method, researchers propose a general learning framework to leverage the latent intents into sequential recommendation via $k$-Means clustering and contrastive learning. Although it has achieved promising performance, it still has two drawbacks as follows.
  1. (1)
    ICLRec adopts a generalized expectation maximization (EM) framework to iteratively encode sequences and perform clustering. In this manner, the representation learning and clustering optimization are separated, leading to sub-optimal performance. In addition, the separated and complex process also increases the difficulty of implementation and reduces reproducibility.
  2. (2)
    The clustering algorithm in ICLRec need to be performed on the full data and can not optimize the clustering distribution via small batch data. Therefore, the scalability of this method is limited, easily leading to out-of-memory or long-running time problems, especially on the large-scale industry dataset.
To alleviate these issues, we propose an end-to-end learnable clustering framework for intent learning in recommendation.
3.4. Proposed Method
To solve the previous problems, we present a new intent learning method termed ELCRec by unifying sequence representation learning into an End-to-end Learnable Clustering framework, for Recommendation. The overall framework of ELCRec is demonstrated in Figure 3. It mainly contains three parts, including sequence encoding, end-to-end learnable cluster module, and cluster-assisted contrastive learning. Next, we introduce them in detail.
3.4.1. Encoding
In this section, we aim to model the users’ behaviors for better recommendation. To be specific, given the user set $\mathcal{U}$, the item set $\mathcal{V}$, and the users’ historical behavior sequence set $\{S^{u}\}_{u=1}^{|\mathcal{U}|}$, the encoder $\mathcal{F}$ embeds the behavior sequences of each user $u$ into the latent space as follows.
(1) $\mathbf{E}^{u}=\mathcal{F}(S^{u}),$
where $\mathbf{E}^{u}\in\mathbb{R}^{|S^{u}|\times d^{\prime}}$ denotes the behavior sequence embedding of user $u$, $d^{\prime}$ is the dimension number of latent features, and $|S^{u}|$ denotes the length of behavior sequence of user $u$. Note that the behavior sequence lengths of different users are different. Therefore, all user behavior sequences are pre-processed to the sequences with the same length $T$ by padding or truncating. Following previous methods (SASRec, ; ICLRec, ), the encoder $\mathcal{F}$ is designed as a Transformer (transformer, ) architecture. Subsequently, to summarize the behaviors of each user, the behavior sequence embedding is aggregated by the concatenate pooling function $\mathcal{P}$ as follows.
(2) $\mathbf{h}_{u}=\mathcal{P}(\mathbf{E}^{u})=\text{concat}(\mathbf{e}^{u}_{1}||...\mathbf{e}^{u}_{i}...||\mathbf{e}^{u}_{T}),$
where $\mathbf{e}^{u}_{i}\in\mathbb{R}^{1\times d^{\prime}}$ denotes the embedding of user behavior at $i$-th step and $\mathbf{h}_{u}\in\mathbb{R}^{1\times Td^{\prime}}$ denotes the aggregated behavior embedding of user $u$. We re-denote $Td^{\prime}$ as $d$ for convenience. By encoding and aggregation, we obtain the behavior embeddings of all users $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$.
3.4.2. End-to-end Learnable Cluster Module
In this section, an End-to-end Learnable Cluster Module (ELCM) is proposed to mine the users’ underlying intents. Concretely, the proposed clustering algorithm groups the users’ behavior embeddings into various clusters, which can represent the latent intents of users.
At first, based on $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$, the cluster centers $\mathbf{C}\in\mathbb{R}^{k\times d}$ are initialized as the learnable neural parameters. Namely, they are the tensors with gradients. Then, we present a clustering loss to train the networks and cluster centers as formulated as follows.
(3) $\mathcal{L}_{\text{cluster}}=\frac{-1}{(k-1)k}\sum_{i=1}^{k}\sum_{j=1,j\neq i}^{k}\left\|\mathbf{c}_{i}-\mathbf{c}_{j}\right\|_{2}^{2}+\\ \frac{1}{bk}\sum_{i=1}^{b}\sum_{j=1}^{k}\left\|\mathbf{h}_{i}-\mathbf{c}_{j}\right\|_{2}^{2},$
where $k$ denotes the number of clusters (intents) and $b$ denotes the batch size. Besides, $\mathbf{h}_{i}\in\mathbb{R}^{1\times d}$ denotes the $i$-th user’s behavior embedding and $\mathbf{c}_{j}\in\mathbb{R}^{1\times d}$ denotes the $j$-th cluster center. In order to better train networks to convergence, we constrain the embeddings to distribute on a unit sphere. Specifically, we apply the $l$-2 normalization to both the user behavior embeddings $\mathbf{H}$ and the cluster centers $\mathbf{C}$ during calculate the clustering loss $\mathcal{L}_{\text{cluster}}$, $\mathbf{H}=\mathbf{H}/\|\mathbf{H}\|_{2}$, $\mathbf{C}=\mathbf{C}/\|\mathbf{C}\|_{2}$.
In Eq. (3), the first term attempts to separate different intents by pushing away different behavior cluster centers. The time complexity and space complexity of the first term are $\mathcal{O}(k^{2}d)$ and $\mathcal{O}(kd)$, respectively, where $k$ is the number of clusters and $d$ is the dimension number of the latent features. Since the number of intents is largely less than the number of users, i.e., $k\ll|\mathcal{U}|$, the first term will not bring significant time or space costs.
In addition, the second term of Eq. (3) aims to pull the behavior embeddings to the intent centers, making the clustering distribution more compact and guiding the network to group similar behaviors into one intent. Here, the behavior embedding $\mathbf{h}_{i}$ is pulled to all center centers $\mathbf{c}_{j},j=1,...,k$ rather than the nearest cluster center. The main reason is that the clustering algorithm is imperfect, and pulling to the nearest center easily leads to the confirmation bias problem (confirmation_bias, ). To this end, the proposed clustering loss optimizes the clustering distribution in an adversarial manner by pulling embeddings together to cluster centers and pushing different cluster centers away. Besides, this term can also optimize the clustering distribution via mini-batch data. Therefore, time complexity and space complexity of the second term are $\mathcal{O}(bkd)$ and $\mathcal{O}(bk+bd+kd)$, respectively. Because the batch size is essentially less than the number of users, namely, $b\ll|\mathcal{U}|$, the second term of clustering loss $\mathcal{L}_{\text{cluster}}$ alleviates the considerable time or space costs.
Differently, in ICLRec (ICLRec, ), the $k$-Means clustering need to be performed on users’ behavior embeddings $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$. In each E-step, it leads to $\mathcal{O}(t|\mathcal{U}|kd)$ time complexity and $\mathcal{O}(|\mathcal{U}|k+|\mathcal{U}|d+kd)$ space complexity, where $t$ denote the iteration steps of $k$-Means. Both the time complexity and space complexity are linear to the number of users, easily leading to the out-of-memory or the long running time problems in the large-scale industrial data. Our proposed end-to-end learnable cluster module can solve these issues. By summarising previous analyses, the overall time and space complexity of calculating the clustering loss $\mathcal{L}_{\text{cluster}}$ are $\mathcal{O}(bkd+k^{2}d+bd)$ and $\mathcal{O}(bk+bd+kd)$, respectively. They are both linear to the batch size $b$. Therefore, it is scalable and plug-and-play, improving the convenience of deployment. For the evidences, please refer to Section 5. Besides, the learnable cluster centers can unify the representation learning and clustering optimization into an end-to-end framework, achieving better recommendation performance. Experimental evidences are listed in Section 4.2 and Section 4.3. In addition, benefiting from the end-to-end framework, the time and space efficiency are both improved, as shown in Section 4.4.
3.4.3. Cluster-assisted Contrastive Learning
After capturing the users’ underlying intents, we aim to collaborate the intent learning and sequential representation learning in this section. Specifically, cluster-assisted contrastive learning (CCL) is proposed to improve the discriminative capability of the users’ behavior.
At first, we conduct contrastive learning among the behavior sequences. Following the general pipeline of contrastive learning (CL4Rec, ; ICLRec, ), we apply the augmentations, including mask, crop, and reorder, to the original user’s behavior. This operation builds two views of behavior sequence of user $u$, i.e., $(S^{u})^{v1}$ and $(S^{u})^{v2}$. According to Section 3.4.1, the behaviors are encoded to the behavior embeddings $\mathbf{h}_{u}^{v1},\mathbf{h}_{u}^{v2}\in\mathbb{R}^{1\times d}$. Then, the sequence contrastive loss of user $u$ is formulated as follows.
(4) $\mathcal{L}_{\text{seq\_cl}}^{u}=-\left(\text{log}\frac{e^{\text{sim}(\mathbf{h}_{u}^{v1},\mathbf{h}_{u}^{v2})}}{\sum_{\text{neg}}e^{\text{sim}(\mathbf{h}_{u}^{v1},\mathbf{h}_{\text{neg}})}}+\text{log}\frac{e^{\text{sim}(\mathbf{h}_{u}^{v1},\mathbf{h}_{u}^{v2})}}{\sum_{\text{neg}}e^{\text{sim}(\mathbf{h}_{u}^{v2},\mathbf{h}_{\text{neg}})}}\right),$
where sim denotes the dot-product similarity, neg denotes the negative samples. Here, the same sequence with different augmentations is recognized as the positive sample pairs, and the other sample pairs are recognized as the negative sample pairs. By minimizing $\mathcal{L}_{\text{seq\_cl}}=\sum_{u}\mathcal{L}_{\text{seq\_cl}}^{u}$, the positive samples are pulled together, and the negative samples are pushed away from each other, therefore enhancing the representation capability of users’ behaviors. During this process, the learned cluster centers are adopted as the self-supervision signals. To be specific, we first query the index of the assigned cluster of the behavior embeddings $\mathbf{h}_{u}^{v1}$ as follows.
(5) $idx=\mathop{\arg\min}\limits_{i}(\|\mathbf{c}_{i}-\mathbf{h}_{u}^{v1}\|_{2}^{2}),$
where $\mathbf{c}_{i}\in\mathbb{R}^{k\times d}$ denotes the $i$-th cluster (intent) center embedding. Then, the intent information is fused to the user behavior during the sequence contrastive learning. Here, we consider two simple fusion strategies, namely, the concatenate fusion $\mathbf{h}_{u}^{v1}=\text{concat}(\mathbf{h}_{u}^{v1}||\mathbf{c}_{idx})$ and the shift fusion $\mathbf{h}_{u}^{v1}=\mathbf{h}_{u}^{v1}+\mathbf{c}_{idx}$. They are optional in our proposed method. The similar operation is apply to the second view of the behavior embedding $\mathbf{h}_{u}^{v2}$. After fusing the intent information to user behaviors, the neural networks are trained by minimizing the sequence contrastive loss $\mathcal{L}_{\text{seq\_cl}}$.
Additionally, to further collaborate intent learning and sequential representation learning, we conduct the contrastive learning between the user’s behaviors and the learnable intent centers. The intent contrastive loss is formulated as follows.
(6) $\mathcal{L}_{\text{intent\_cl}}^{u}=-\left(\text{log}\frac{\min_{i}e^{\text{sim}(\mathbf{h}_{u}^{v1},\mathbf{c}_{i})}}{\sum_{\text{neg}}e^{\text{sim}(\mathbf{h}_{u}^{v1},\mathbf{c}_{\text{neg}})}}+\text{log}\frac{\min_{i}e^{\text{sim}(\mathbf{h}_{u}^{v2},\mathbf{c}_{i})}}{\sum_{\text{neg}}e^{\text{sim}(\mathbf{h}_{u}^{v2},\mathbf{c}_{\text{neg}})}}\right),$
where $\mathbf{h}_{u}^{v_{1}},\mathbf{h}_{u}^{v_{2}}$ are two-view behavior embedding of the user $u$. Besides, neg denotes the negative behavior-intent pairs among all pairs. Here, we regard the behavior embedding and the corresponding nearest intent center as the positive pair, and others are regarded as negative pairs. By minimizing the intent contrastive loss $\mathcal{L}_{\text{intent\_cl}}=\sum_{u}\mathcal{L}_{\text{intent\_cl}}^{u}$, behaviors with the same intents are pulled together, but behaviors with different intents are pushed away.
In this section, the learn clusters are guided to assist the contrastive learning for sequential representation learning. In summary, the objective of cluster-assisted contrastive learning is formulated as $\mathcal{L}_{\text{ccl}}=\mathcal{L}_{\text{seq\_cl}}+\mathcal{L}_{\text{intent\_cl}}$. In this manner, the clustering optimization and representation learning are unified into an end-to-end framework, improving the recommendation performance and convenience of deployment. The effectiveness of CCL is verified in Section 4.3.
3.4.4. Overall Objective
The neural networks and learnable clusters are trained with multiple tasks, including clustering, cluster-assisted contrastive learning, and next-item prediction. The clustering task aims to capture the users’ underlying intents. Besides, the cluster-assisted contrastive learning aims to collaborate the intent learning and sequential representation learning. In addition, the next-item prediction task is a widely used task for recommendation system (SASRec, ; S3_rec, ; ComiRec, ; ICLRec, ). In summary, the overall objective of the proposed ELCRec is formulated as follows.
(7) $\mathcal{L}_{\text{overall}}=\mathcal{L}_{\text{next\_item}}+0.1\times\mathcal{L}_{\text{ccl}}+\alpha\times\mathcal{L}_{\text{cluster}},$
where $\mathcal{L}_{\text{next\_item}}$, $\mathcal{L}_{\text{ccl}}$, and $\mathcal{L}_{\text{cluster}}$ denotes the next item prediction loss, cluster-assisted contrastive learning loss, and clustering loss, respectively. Besides, $\alpha$ is a trade-off hyper-parameter.
4. Experiment
This section aims to comprehensively evaluate our proposed ELCRec by answering the following research questions.
  1. (1)
    Superiority: does ELCRec outperform the existing state-of-the-art sequential recommendation methods?
  2. (2)
    Effectiveness: are the proposed end-to-end learnable cluster module and cluster-assisted contrastive learning effective?
  3. (3)
    Efficiency: how about the time and memory efficiency of the proposed ELCRec?
  4. (4)
    Sensitivity: what is the performance of the proposed method with different hyper-parameters?
  5. (5)
    Convergence: will the proposed loss function and the recommendation performance converge well?
4.1. Setup
4.1.1. Environment
Experimental results on the public benchmarks are obtained from the desktop computer with one NVIDIA GeForce RTX 4090 GPU, six 13th Gen Intel(R) Core(TM) i9-13900F CPUs, and the PyTorch platform.
4.1.2. Dataset
We performed our experiments on four public datasets: Sports, Beauty, Toys, and Yelp [^1] https://www.yelp.com/dataset. The Sports, Beauty, and Toys datasets are subcategories of the Amazon Review Dataset (dataset, ). The Sports dataset contains reviews for sporting goods, the Beauty dataset contains reviews for beauty products, and the Toys dataset contains reviews for toys. The Yelp dataset, on the other hand, focuses on business recommendations and is provided by Yelp company. Table 7 provides a summary of the datasets’ details. We followed the approach of previous methods (ICLRec, ; S3_rec, ; CL4Rec, ) and only kept datasets where all users and items have at least five interactions. Additionally, we adopted the dataset split settings used in the previous method (ICLRec, ).
4.1.3. Metric
To comprehensively evaluate our proposed method, we adopt two groups of metrics, including Hit Ratio @ $k$ (HR@$k$), and Normalized Discounted Cumulative Gain @ $k$ (NDCG@$k$), where $k\in\{5,20\}$.
4.1.4. Baseline
We compare our method with nine baselines including BPR-MF (BPR_MF, ), GRU4Rec (GRU4Rec, ), Caser (Caser, ), SASRec (SASRec, ), DSSRec (DSSRec, ), BERT4Rec (BERT4Rec, ), S3-Rec (S3_rec, ), CL4SRec (CL4Rec, ), and ICLRec (ICLRec, ). For a detailed introduction to these methods, please refer to related work in Section 2.2.
4.1.5. Implementation
For the baselines, we utilized the implementation provided in (ICLRec, ). Our proposed method, ELCRec, was implemented using the PyTorch deep learning platform. In the Transformer encoder, we employed self-attention blocks with 2 attention heads. The latent dimension, denoted as $d$, was set to 64, and the maximum sequence length, denoted as $T$, was set to 50. We utilized the Adam optimizer with a learning rate of 1e-03. The decay rate for the first moment estimate was set to 0.9, and the decay rate for the second moment estimate was set to 0.999. The cluster number, denoted as $k$, was set to 256 for the Yelp and Beauty datasets, and 512 for the Sports and Toys datasets. The trade-off hyper-parameter, denoted as $\alpha$, was set to 1 for the Sports and Toys datasets, 0.1 for the Yelp dataset, and 10 for the Beauty dataset. During training, we monitored the training process using the Weights & Biases platform.
4.2. Superiority
In this section, we aim to answer research question (1) and demonstrate the superiority of the proposed ELCRec method. To be specific, we compare ELCRec with nine state-of-the-art recommendation baselines (BPR_MF, ; GRU4Rec, ; Caser, ; SASRec, ; DSSRec, ; BERT4Rec, ; S3_rec, ; CL4Rec, ; ICLRec, ). Experimental results are the mean values of three runs. As shown in Table 1, the bold and underlined values denote the best and runner-up results, respectively. From these results, we have three conclusions as follows.
  1. (1)
    Firstly, the non-sequential model BPR-MF (BPR_MF, ) can not achieve promising recommendation performance since the shallow method lacks the representation learning capability of users’ historical behaviors.
  2. (2)
    Besides, the standard sequential methods (GRU4Rec, ; Caser, ; SASRec, ) improve the recommendation via different deep neural networks such as CNN (CNN, ), RNN (RNN, ), and Transformer (transformer, ). However, they can not beat our method because the self-supervision signal is limited.
  3. (3)
    Additionally, the recent methods (BERT4Rec, ; S3_rec, ; CL4Rec, ) enhance the self-supervised capability of recommendation systems via the self-supervised learning techniques. However, they neglect the underlying users’ intent, thus leading to sub-optimal performance.
  4. (4)
    More recently, the intent learning methods (DSSRec, ; ICLRec, ) have been proposed to mine users’ underlying intent to assist recommendation. Motivated by their success, we propose an end-to-end learnable cluster framework named ELCRec for intent learning. Befitting from the strong intent mining capability of ELCRec, it surpasses other intent learning methods. To further verify the superiority of ELCRec, we conduct the $t$-test between the best and runner-up methods. As shown in Figure 1, the most $p$-value is less than 0.05 except HR@5 on the Toys dataset. This indicates our proposed ELCRec significantly outperforms the runner-up method.
Overall, the extensive experiments demonstrate the superiority of our proposed method.
Table 1. Recommendation performance on public benchmarks. The bold and underlined values denote the best and runner-up results, respectively. $*{}^{*}$ indicates that, in the $t$-test, the best method significantly outperforms the runner-up with $p<0.05$.
Dataset Metric BPR-MF GRU4Rec Caser SASRec DSSRec BERT4Rec S3-Rec CL4SRec ICLRec ELCRec $p$-value
(BPR_MF, ) (GRU4Rec, ) (Caser, ) (SASRec, ) (DSSRec, ) (BERT4Rec, ) (S3_rec, ) (CL4Rec, ) (ICLRec, ) Ours
Sports HR@5 0.0141 0.0162 0.0154 0.0206 0.0214 0.0217 0.0121 0.0217 0.0263 0.0286 $*{}^{*}$
HR@20 0.0323 0.0421 0.0399 0.0497 0.0495 0.0604 0.0344 0.0540 0.0630 0.0648 $*{}^{*}$
NDCG@5 0.0091 0.0103 0.0114 0.0135 0.0142 0.0143 0.0084 0.0137 0.0173 0.0185 $*{}^{*}$
NDCG@20 0.0142 0.0186 0.0178 0.0216 0.0220 0.0251 0.0146 0.0227 0.0276 0.0286 $*{}^{*}$
Beauty HR@5 0.0212 0.0111 0.0251 0.0374 0.0410 0.0360 0.0189 0.0423 0.0495 0.0529 $*{}^{*}$
HR@20 0.0589 0.0478 0.0643 0.0901 0.0914 0.0984 0.0487 0.0994 0.1072 0.1079 $*{}^{*}$
NDCG@5 0.0130 0.0058 0.0145 0.0241 0.0261 0.0216 0.0115 0.0281 0.0326 0.0355 $*{}^{*}$
NDCG@20 0.0236 0.0104 0.0298 0.0387 0.0403 0.0391 0.0198 0.0441 0.0491 0.0509 $*{}^{*}$
Toys HR@5 0.0120 0.0097 0.0166 0.0463 0.0502 0.0274 0.0143 0.0526 0.0586 0.0585 1.22E-01
HR@20 0.0312 0.0301 0.0420 0.0941 0.0975 0.0688 0.0235 0.1038 0.1130 0.1138 $*{}^{*}$
NDCG@5 0.0082 0.0059 0.0107 0.0306 0.0337 0.0174 0.0123 0.0362 0.0397 0.0403 $*{}^{*}$
NDCG@20 0.0136 0.0116 0.0179 0.0441 0.0471 0.0291 0.0162 0.0506 0.0550 0.0560 $*{}^{*}$
Yelp HR@5 0.0127 0.0152 0.0142 0.0160 0.0171 0.0196 0.0101 0.0229 0.0233 0.0236 $*{}^{*}$
HR@20 0.0346 0.0371 0.0406 0.0443 0.0464 0.0564 0.0314 0.0630 0.0645 0.0653 $*{}^{*}$
NDCG@5 0.0082 0.0091 0.0080 0.0101 0.0112 0.0121 0.0068 0.0144 0.0146 0.0150 $*{}^{*}$
NDCG@20 0.0143 0.0145 0.0156 0.0179 0.0193 0.0223 0.0127 0.0256 0.0261 0.0266 $*{}^{*}$
4.3. Effectiveness
This section is dedicated to addressing research question (2) and evaluating the effectiveness of the End-to-end Learnable Cluster Module (ELCM) and Cluster-assisted Contrastive Learning (CCL). To achieve this, we conducted meticulous ablation studies on four publicly available datasets. Figure 4 illustrates the results, where sub-figures (a), (b), (c), and (d) represent the Sports, Beauty, Toys, and Yelp datasets, respectively. The results are aggregated from four metrics: HR@5, HR@20, NDCG@5, and NDCG@20. In each sub-figure, the labels “B,” “B+CCL,” “B+ELCM,” and “ELCRec” correspond to the baseline, baseline with CCL, baseline with ELCM, and baseline with both CCL and ELCM, respectively. Through the ablation studies, we draw three key conclusions.
Refer to caption (a) Sports
Refer to caption (b) Beauty
Refer to caption (c) Toys
Refer to caption (d) Yelp
Figure 4. Ablation studies of the proposed end-to-end learnable cluster module and the cluster-assisted contrastive learning. The results are the sum of four metrics, including HR@5, HR@20, NDCG@5, and NDCG@20.
  1. (1)
    Firstly, “B+CCL” can outperform the baseline on two out of four datasets. Concretely, it indicates that the proposed cluster-assisted contrastive learning is effective on the Sports and Toys datasets. However, the cluster centers are not optimized well via $k$-Means clustering, thus leading to sub-optimal performance on the Beauty and Yelp datasets.
  2. (2)
    Secondly, we observe that “B+ELCM” can surpass the baseline “B” significantly on all benchmarks. This demonstrates that our proposed end-to-end learnable cluster module helps the model better capture the users’ underlying intents, therefore achieving promising performance.
  3. (3)
    Thirdly, the baseline with CCL and ELCM (i.e., “ELCRec”) achieves the best performance on three out of four datasets. It verifies the effectiveness of the combination of these two modules. On the Toys dataset, “ELCRec” can outperform the baseline but achieve worse performance than “B+CCL” and “B+ELCM”. This phenomenon indicates it is worth doing research on the better collaboration of these two modules.
To summarize, the extensive ablation studies verify the effectiveness of modules in the proposed ELCRec.
4.4. Efficiency
In this section, we test the efficiency of ELCRec and answer the research question (3). Specifically, the efficiency contains two perspectives, including training time costs and GPU memory costs. We summarize two observations as follows.
  1. (1)
    The results of training time costs are listed in Table 2. We find that our proposed ELCRec can speed up ICLRec on three out of four datasets. Overall, on four datasets, the training time is decreased by 7.18% on average.
  2. (2)
    The results of GPU memory costs are listed in Table 3. The results demonstrate that the GPU memory costs of our ELCRec are lower than that of ICLRec on four datasets. On average, the GPU memory costs are decreased by 9.58%.
In summary, both the time efficiency and memory efficiency of our proposed ELCRec are verified on four benchmarks.
Table 2. Time costs (in second) on four public benchmarks.
Dataset Sports Beauty Toys Yalp Avg.
ICLRec 5282 3770 4374 4412 4460
ELCRec 5360 2922 4124 4151 4139
Ratio 1.48% $\uparrow$ 22.49% $\downarrow$ 5.72% $\downarrow$ 5.92% $\downarrow$ 7.18% $\downarrow$
Table 3. GPU memory costs (in MB) on public benchmarks.
Dataset Sports Beauty Toys Yelp Avg.
ICLRec 1944 1798 2887 3671 2575
ELCRec 1781 1594 2555 3383 2328
Ratio 8.38% $\downarrow$ 11.35% $\downarrow$ 11.50% $\downarrow$ 7.85% $\downarrow$ 9.58% $\downarrow$
5. Application
Benefiting from the versatility of the proposed ELCRec, we aim to apply it to the real-world industrial scenario. The applied scenario is the live-streaming recommendation on the front page of the Alipay app. The UV and PV of this application is about 50 million and 130 million, respectively. Most users are new to this application, therefore leading to the sparsity of users’ behaviors. To solve this cold-start problem in the recommendation system, we adopt the proposed end-to-end learnable cluster module to group users and recommend. Firstly, due to the sparsity of users’ behaviors, we replace the users’ behavior with the users’ activities and model them via the multi-gate mixture-of-expert (MMOE) (MMOE, ). Then, the end-to-end learnable cluster module is adopted to group the users into various clusters. Through this module, the high-activity users and new users are grouped into different clusters, alleviating the cold-start issue and assisting in better recommendations. Eventually, the networks are trained with the multi-task training strategy. Next, we introduce the results and case studies on the industrial data.
5.1. Industrial Result
We test the online recommendation performance of our proposed module. Concretely, the overall recommendation performance is listed in Table 4. The metrics contain two parts, including live-streaming metrics (PVCTR and VV) and commercial metrics (PVCTR and UVCTR). “B” denotes the baseline and “B+ELCM” denotes the baseline with the proposed end-to-end learnable cluster module. The results indicate that our proposed module can improve the recommendation performance of the baseline by about 2%. Besides, “B+ELCM” outperforms “B” significantly with $p<0.05$ with three out of four metrics.
Table 4. Experimental results of the online recommendation system. The bold values denotes the significant improvements with $p<0.05$.
Method Live-streaming Metric Commercial Metric
PVCTR VV PVCTR UVCTR
B - - - -
B+ELCM 2.45% $\uparrow$ 2.28% $\uparrow$ 2.41% $\uparrow$ 1.62% $\uparrow$
In addition, we further analyze the recommendation performance on different user groups, as shown in Table 5. Based on the user activity, we classify them into five groups, including Pure New users (PN), New users (N), Low Activity users (LA), Medium Activity users (MA), and High Activity users (HA). Compared with the general recommendation algorithm that is unfriendly to new users, the experimental results show that our module not only improves the recommendation performance of high activity users but also improves the recommendation performance of new users. Therefore, it can alleviate the cold-start problem and construct a more friendly user ecology.
Table 5. Experimental results of different user groups in the online recommendation system.
Metric PN N LA MA HA
PVCTR 6.96% $\uparrow$ 1.67% $\uparrow$ 1.98% $\uparrow$ 0.35% $\uparrow$ 19.02% $\uparrow$
VV 6.81% $\uparrow$ 1.50% $\uparrow$ 1.50% $\uparrow$ 0.04% $\uparrow$ 16.90% $\uparrow$
5.2. Case Study
In this section, we conduct case studies on the user group distribution and the learned cluster centers. Firstly, for the user group distribution, the results are demonstrated in Figure 8. We visualize the cluster distribution of different user groups. “top” denotes the cluster IDs that have the highest proportion in the user group. “bottom” denotes the cluster IDs that have the lowest proportion in the user group. Two findings are listed as follows. (1) When the user activity increases, the distribution becomes sharper. Namely, most of this group of users tend to distribute into one or two clusters. (2) The “top” cluster IDs of the high activity user group, such as cluster 10 and cluster 8, are exactly the “bottom” cluster IDs of the low activity user group. Similarly, the “bottom” cluster IDs of the high activity user group, such as cluster 9, are exactly the “top” cluster IDs of the low activity user group. This indicates that the learned cluster centers can separate different user groups.
Secondly, we also conduct extensive case studies on the learned clusters. To be specific, we analyze the user distribution of each cluster, as shown in Figure 9. From the results, we can observe that, in cluster 6, most users are new. Besides, in the cluster 8, the most users are with medium activity. In addition, in cluster 10, most users are with high activity and medium activity. Moreover, in cluster 13, most users are with low activity and medium activity. Previous observations show that the learned centers can separate the users into different groups.
In summary, these case studies further verify the effectiveness of our proposed method. Also, they provide various insights for future work.
6. Conclusion
The topic of this paper is intent learning in recommendation systems. Concretely, the intent learning methods aim to mine the users’ underlying intents and perform recommendations based on their learned intents. In this paper, we carefully analyze the drawbacks of previous methods. The complex and cumbersome alternating optimization limits the scalability and performance of the existing method. To this end, we propose an end-to-end learnable clustering framework named ELCRec for intent learning in recommendation systems. To be specific, the cluster centers are initialized as the learnable neural parameters. Then, the proposed clustering loss optimizes the cluster distribution based on the mini-batch data, enabling the scalability of ELCRec. Besides, to further improve the recommendation performance, the learned cluster centers are assigned as the self-supervision signals for sequential representation learning. Extensive experiments demonstrate ELCRec’s six abilities, including superiority, effectiveness, efficiency, sensitivity, and convergence. In addition, benefiting from the versatility of ELCRec, we successfully apply it to the real-world industrial scenario and also achieve promising performance. In the future, it is worth developing the cluster number-free clustering framework for intent learning based on the sample density.
Appendix A Appendix
A.1. Notation & Dataset
The basic notations are summarized in Table 6. The statistical information of four public benchmarks are summarized in Table 7.
Table 6. Basic notations.
Notation Meaning
$\mathcal{U}$ User set
$\mathcal{V}$ Item set
$\{S^{u}\}_{u=1}^{|\mathcal{U}|}$ Users’ behavior sequence set
$d^{\prime}$ Dimension number of latent features
$d$ Dimension number of aggregated latent features
$b$ Batch size
$k$ Cluster number
$T$ Maximum sequence length
$\mathcal{L}_{\text{cluster}}$ Clustering loss
$\mathcal{L}_{\text{seq\_cl}}$ Behavior sequence contrastive loss
$\mathcal{L}_{\text{intent\_cl}}$ Intent contrastive loss
$\mathcal{L}_{\text{ccl}}$ Cluster-assisted contrastive learning loss
$\mathcal{L}_{\text{next\_item}}$ Next item prediction loss
$\mathcal{L}_{\text{overall}}$ Overall loss of the proposed ELCRec
$\mathcal{F}$ Encoder network
$\mathcal{P}$ Concatenate pooling function
$\mathbf{E}^{u}\in\mathbb{R}^{|S^{u}|\times d^{\prime}}$ Behavior sequence embedding of user $u$
$\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$ Behavior embeddings of all users
$\mathbf{H}^{v_{k}}\in\mathbb{R}^{|\mathcal{U}|\times d}$ Behavior embeddings of all users in view $v_{k}$
$\mathbf{C}\in\mathbb{R}^{k\times d}$ Learnable cluster center embeddings
Table 7. Statistical information of four public datasets.
Dataset #User #Item #Action Avg. Len. Sparsity
Sports 35,598 18,357 0.3M 8.3 99.95%
Beauty 22,363 12,101 0.2M 8.9 99.95%
Toys 19,412 11,924 0.17M 8.6 99.93%
Yelp 30,431 20,033 0.3M 8.3 99.95%
A.2. Algorithm
Due to the limitation of the pages, we present the overall algorithm of the proposed ELCRec method in Algorithm 1.
Algorithm 1 End-to-end Learnable Clustering Framework for Recommendation (ELCRec)
Input: user set $\mathcal{U}$; item set $\mathcal{V}$; historical behavior sequences $\{\mathcal{S}^{u}\}_{u=1}^{|\mathcal{U}|}$; cluster number $k$; epoch number $E$; learning rate; trade-off parameter $\alpha$.
Output: Well-trained recommendation model.
1:  Initialize model parameters in encoders.
2:  for $\text{epoch}=1,2,...,E$ do
3:     # Encoding
4:     for $\text{u}=1,2,...,|\mathcal{U}|$ do
5:        Obtain $u$-th user’s behavior sequence embedding $\mathbf{E}^{u}\in\mathbb{R}^{|\mathcal{S}^{u}|\times d^{\prime}}$ via encoding $\mathcal{S}^{u}$ in Eq. (1).
6:        Obtain $u$-th user’s aggregated behavior embedding $\mathbf{h}_{u}\in\mathbb{R}^{1\times d}$ via aggregating $\mathbf{E}^{u}$ in Eq. (2)
7:     end for
8:     Obtain behavior embeddings of all users $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$.
9:     # End-to-end Learnable Cluster Module
10:     Initialize cluster center embeddings $\mathbf{C}\in\mathbb{R}^{k\times d}$ as tensors with gradients.
11:     Calculating clustering loss to train networks and cluster centers in Eq. (3).
12:     Apply $l$-2 normalization to both $\mathbf{H}$ and $\mathbf{C}$ during training.
13:     # Cluster-assisted Contrastive Learning
14:     Apply data augmentations to generate two views of behavior sequences.
15:     Encode the two views of the behavior sequences and obtain the embeddings.
16:     Calculating $\mathcal{L}_{\text{seq\_cl}}$ to conduct contrastive learning among behavior sequences.
17:     Query the index of assigned cluster of the behavior embeddings via Eq. (5).
18:     Fuse the intent information to behavior embeddings via concatenate or shift.
19:     Calculating $\mathcal{L}_{\text{intent\_cl}}$ to conduct intent contrastive learning in Eq. (6).
20:     Calculating $\mathcal{L}_{\text{next\_item}}$ to conduct next item prediction task.
21:     Minimizing $\mathcal{L}_{\text{overall}}$ to train networks and cluster centers in Eq. (7).
22:  end for
23:  Return Trained ELCRec model.
A.3. Deployment Details
We aim to apply our proposed method to the real-world industrial scenario. Concretely, the ELCRec algorithm is applied to the front page of the Alipay app. The user view (UV) and page view (PV) of this application are about 50 million and 130 million, respectively. Since most of the users are new to this application, it easily leads to the sparsity of users’ behaviors, namely, the cold-start problem in recommendation systems. Our proposed ELCRec can alleviate this problem by grouping users and then making recommendations. Firstly, we introduce the online baseline of this project. Since the sparsity of the users’ behaviors, we replaced the users’ behaviors with the users’ activities. Then, the online baseline multi-gate mixture-of-expert (MMOE) (MMOE, ) models the users’ activities. In this model, the experts are designed to extract the features of users, and the multi-gates are designed to select specific experts. The inputs of the multi-gates are the activities of the users. This design aims to train an activity-awarded model to group different users and then conduct recommendations. However, we found the performance of this model is limited, and the output of the gates is smooth, indicating that this model may fail to group users. Meanwhile, on the open benchmarks, extensive experiments demonstrate the proposed end-to-end learnable clustering module is effective and scalable. Thus, to solve the above problem, ELCRec is adopted in this project. It is designed to assist the gate to group users. For example, the high-activity users and new users are grouped into different clusters, and then the users in different clusters will be recommended differently. Therefore, it alleviates the cold-start issue and further improves the recommendation performance. It is worth mentioning that the networks are trained with multi-task targets, e.g., CTR, CVR, etc. Following the previous online baseline, the method is implemented with the TensorFlow deep learning platform (tensorflow, ).
A.4. Additional Test on Diverse Domains
We acknowledge that the applied domains of our proposed method are limited. We aim to test our proposed ELCRec on other domains, such as news recommendation and movie recommendation.
A.4.1. Movie Recommendation
Specifically, for movie recommendation, we conducted experiments on the MovieLens 1M dataset (ML-1M). This dataset contains 1 million ratings from 6000 users on 4000 movies. In this experiment, we compared our proposed ELCRec with the baseline ICLRec. The experimental results are presented in the table below.
Table 8. Values marked with a star (*) indicate that the $p$-value is less than 0.05. Bold value denotes the best value.
Method HR@5 HR@20 NDCG@5 NDCG@20
ICLRec 0.0293 0.0777 0.0186 0.0320
ELCRec 0.0333 0.0836 0.0208 0.0347
$p$-value 4.03E-06 6.68E-09 6.36E-06 1.66E-06
From these experimental results, we can draw three conclusions. (1) Our proposed ELCRec achieves better recommendation performance, as evidenced by higher values for all four metrics: HR@5, HR@20, NDCG@5, and NDCG@20. For example, with the HR@5 metric, ELCRec outperforms ICLRec by 13.56%. (2) We calculated the $p$-value between our method and the baseline ICLRec. The results indicate that all the $p$-values are less than 0.05, suggesting that our ELCRec significantly outperforms ICLRec.
A.4.2. News Recommendation
Additionally, for news recommendation, we plan to conduct experiments on the MIND dataset. This dataset is a large-scale dataset for news recommendation research, collected from anonymized behavior logs of the Microsoft News website. The MIND dataset contains about 160k English news articles and more than 15 million impression logs generated by 1 million users. Due to limitations in computational resources and the scalability of this dataset, we may not be able to provide the experimental results during the rebuttal period. However, we assure you that we will update the paper with the experimental results once they are obtained.
A.5. Sensitivity
This section aims to answer research question (4). To verify the sensitivity of the proposed ELCRec, we test the performance on four datasets with different hyper-parameters. The experimental results are demonstrated in Figure 5. The x-axis denotes the values of hyper-parameters, and the y-axis denotes the values of the HR@5 metric. Concretely, for the trade hyper-parameter $\alpha$, we test the performance with $\alpha\in\{0.01,0.1,1,10,100\}$. We find that our proposed ELCRec is not very sensitive to trade-off $\alpha$. And ELCRec can achieve promising performance when $\alpha\in[0.1,10]$. Besides, for the cluster number $k$, we test the recommendation performance with $\alpha\in\{32,64,128,256,512\}$. The results show that ELCRec is also not very sensitive to cluster number $k$ and can perform well when $k\in[256,512]$.
Refer to caption (a) Trade-off $\alpha$
Refer to caption (b) Cluster number $k$
Figure 5. Sensitivity analyses of ELCRec. The results are evaluated by the HR@5 metric.
Refer to caption (a) Sports
Refer to caption (b) Beauty
Refer to caption (c) Toys
Refer to caption (d) Yelp
Figure 6. Convergence analyses. The first and second row denotes HR@5 on the evaluation set and training loss, respectively.
A.6. Convergence
To answer the research question (5), we monitor the recommendation performance and training loss as shown in Figure 6. We find that the losses gradually decrease and eventually converge. Besides, during the training process, the recommendation performance gradually increases and eventually reaches a promising value.
A.7. Visualization
To further demonstrate ELCRec’s capability to capture users’ underlying intents, we conduct visualization experiments on four public datasets. Concretely, the learned behavior sequence embeddings are visualized via $t$-SNE during the training process. As shown in Figure 7, the first row to the fourth row denotes the results on Sports, Beauty, Toys, and Yelp, respectively. From these experimental results, we have three observations as follows. (1) At the beginning of training, the behavior sequence embeddings are disorganized and can not reveal the underlying intent of users. (2) During the training process, the latent distribution is optimized, and similar behaviors are grouped into the underlying intent. (3) After training, the users’ underlying intents appear, and we highlight them with circles in Figure 7. These intents can assist recommendation systems in better modeling users’ behavior and recommending items. In summary, the above experiments and observations verify the effectiveness of our proposed ELCRec.
A.8. Case Study
The results of case studies are demonstrated in Figure 8 and Figure 9. The detailed analyses refer to Section 5.2.
Refer to caption Refer to caption Refer to caption Refer to caption (a) 0 epoch
Refer to caption Refer to caption Refer to caption Refer to caption (b) 50 epoch
Refer to caption Refer to caption Refer to caption Refer to caption (c) 100 epoch
Refer to caption Refer to caption Refer to caption Refer to caption (d) 150 epoch
Refer to caption Refer to caption Refer to caption Refer to caption (e) 200 epoch
Refer to caption Refer to caption Refer to caption Refer to caption (f) final epoch
Figure 7. t-SNE visualization on four public datasets.
Refer to caption (a) New users
Refer to caption (b) Low activity users
Refer to caption (c) Medium activity users
Refer to caption (d) High activity users
Figure 8. Case studies on different user groups. The distributions of different user groups are visualized. “top” denotes the cluster IDs, which have the highest proportion in the user group. “bottom” denotes the cluster IDs, which have the lowest proportion in the user group.
Refer to caption (a) Cluster 6
Refer to caption (b) Cluster 8
Refer to caption (c) Cluster 10
Refer to caption (d) Cluster 13
Figure 9. Case studies on the learned cluster centers. “HA”, “MA”, “LA”, and “N” denotes the high activity, medium activity, low activity, and new user groups.
A.9. Solutions to Balance Problem
Balancing the different loss functions in our method is indeed an important challenge. Our overall loss function consists of next-item prediction loss, cluster-assisted contrastive loss, and cluster loss. It is formulated as follows: $\mathcal{L}_{\text{overall}}=\mathcal{L}_{\text{next\_item}}+0.1\times\mathcal{L}_{\text{ccl}}+\alpha\times\mathcal{L}_{\text{cluster}}$. We set the weight of $\mathcal{L}_{\text{ccl}}$ as 0.1 to maintain it in the same order of magnitude as the first term. This reduces the number of hyper-parameters and simplifies the selection process. The weight of $\mathcal{L}_{\text{cluster}}$ is set as a hyper-parameter $\alpha$. We test different values of $\alpha\in\{0.01,0.1,1,10,100\}$ and find that our ELCRec method is not very sensitive to the trade-off $\alpha$. Promising performance is achieved when $\alpha\in[0.1,10]$. The sensitivity analysis experiments are presented in Figure 5 (b) of the Appendix. In our proposed model, we set $\alpha$ to 1 for Sports and Toys datasets, 0.1 for the Yelp dataset, and 10 for the Beauty dataset. The selection of $\alpha$ is mainly based on the model performance. Balancing trade-offs in models with multiple loss functions or tasks is a challenge in many deep learning methods. There are various strategies to address this issue.
  • Weighted Balancing. Assign weights to each loss function to control their contribution to the overall loss. By adjusting the weights, a balance can be achieved between different loss functions. This can be determined through prior knowledge, empirical rules, or methods like cross-validation. (our method)
  • Dynamic Weight Adjustment. Adjust the weights of the loss functions in real-time based on the model’s training progress or the characteristics of the data. For example, dynamically adjust the weights based on the model’s performance on a validation set, giving relatively smaller weights to underperforming loss functions.
  • Multi-objective Optimization. Treat different loss functions as multiple optimization objectives and use multi-objective optimization algorithms to balance these objectives. This allows for the simultaneous optimization of multiple loss functions and seeks balance between them.
  • Gradient-based Adaptive Adjustment. Adaptively adjust the weights of loss functions based on their gradients. If a loss function has a larger gradient, it may have a greater impact on the model’s training, and its weight can be increased accordingly.
  • Ensemble Methods. Train multiple models based on different loss functions and use ensemble learning techniques to combine their prediction results. By combining the predictions of different models, a balance between different loss functions can be achieved.
In the future, we will continue to improve our model by refining the weighting between different loss functions based the above strategies.
A.10. Theoretical Analyses
In this subsection, we investigate the generalization bounds of the proposed clustering loss. Our analysis is based on the Rademacher complexity and investigates how the proposed clustering loss improves the generalization bound of the algorithm.
Without loss of generality, we have the following notation. Let $\mathbf{x}\in\mathcal{X}$ be the input, where $\mathbf{x}$ are generated from a underlying distribution $\mathbf{x}\sim\mathcal{P}$. Given $n$ training samples $\mathcal{S}\triangleq\{\boldsymbol{x}_{i}\}_{i\in[n]}$ generated from distribution $\mathcal{P}$, we denote its empirical distribution by $\mathcal{P}^{n}$. For every hyperparameter $\omega\in\Omega$, we define $\mathcal{F}_{\omega}$ as a distribution-dependent hypothesis space corresponding to the $\omega$, where $\Omega$ is a finite set of hyperparameters. $\mathcal{F}_{\omega}$ is defined as $\{f_{\omega}|f_{\omega}=\mathcal{A}_{\omega}(S),S\in\mathcal{S}\}$, where $\mathcal{A}_{\omega}$ is an algorithm that outputs the hypothesis $f_{\omega}$ given a dataset $S$.
In the subsequent analysis, we denote $\mathcal{L}_{\text{cluster}}(S,f_{\omega})=\ell\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)$ as the proposed cluster loss $\mathcal{L}_{\text{cluster}}$ with the embedding $\mathbf{c}$. Let $u,v$ are the upper and lower bounds of the cluster loss respectively. In other words, $u\geq\ell\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)\geq v$. In this paper, $u=4$ and $v=-4$. $\mathcal{R}^{\ell}_{n}{(\mathcal{F}_{\omega})}$ is the rademacher complexity of the set $\left\{\mathbf{x}\mapsto\ell(f_{\omega}(\mathbf{x},\mathbf{c}):f_{\omega}\in\mathcal{F}_{\omega}\right\}$. Besides, we have $\mathbb{E}_{\mathbf{x}\sim\mathcal{P}^{n}}\left[\ell(f_{\omega}(\mathbf{x},\mathbf{c})\right]=\frac{1}{n}\sum_{i=1}^{n}\ell\left(f_{\omega}(\mathbf{x}_{i},\mathbf{c})\right)$.
With the notation above, we have the following theorem
Theorem 1 ().
For any $\delta>0$ and $\omega\in\Omega$, for all $f_{\omega}\in\mathcal{F}_{\omega}$, with the probability at least $1-\delta$, we have:
(8) $\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}[\ell(f_{\omega}(\mathbf{x},\mathbf{c}))]-\mathbb{E}_{\mathbf{x}\sim\mathcal{P}^{n}}[\ell(f_{\omega}(\mathbf{x},\mathbf{c}))]$
[…]
where $\text{ln}\Pi_{\mathcal{F_{\omega}}}(n)$ denotes the growth function.
Remark 1 ().
[…]
To prove the above theorem, we need the following lemma
Lemma 2 ().
(lemma_1, ) Let $\mathcal{F}$ be a class of real-valued function that map from $\mathcal{X}$ to $[v,u]$. Let $\mathcal{D}$ be a probability distribution on $\mathcal{X}\times[v,u]$, and suppose that sample set $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$ are chosen independently according to the distribution $\mathcal{D}$. For all $f\in\mathcal{F}$, with probability at least $1-\delta$, we have:
[…]
where $\Phi(S)=\sup_{f\in\mathcal{F}}\left(\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}[f]-\mathbb{E}_{\mathbf{x}\sim\mathcal{P}^{n}}[f]\right)$, $\mathcal{R}_{n}(\cdot))$ is the correspondent rademacher complexity.
Lemma 3 ().
(lemma_2, ) Let $\mathcal{F}$ be the hypothesis space. The Rademacher complexity $\mathcal{R}_{n}(\mathcal{F})$ and the growth function $\Pi_{\mathcal{F}}(n)$ have:
[…]
Proof.
With the above lemma, we have the following derivation
(11) $Let\ \Phi(S)$ $=\sup_{f_{\omega}\in\mathcal{F}_{\omega}}\left(\mathbb{E}_{\mathbf{x}\sim P}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)\right]-\mathbb{E}_{\mathbf{x}\sim P^{n}}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)\right]\right)$
$=\sup\left(\mathbb{E}_{\mathbf{x}\sim P}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)\right]-\frac{1}{n}\sum_{i=1}^{n}\left[\mathcal{L}\left(f_{\omega}(x_{i},\mathbf{c})\right)\right]\right).$
We first provide an upper bound on $\Phi(S)$ by using McDiarmid’s inequality. To apply McDiarmid’s inequality, we compute an upper bound on $\left|\Phi(S)-\Phi(S^{\prime})\right|$ where $S$ and $S^{\prime}$ be two training datasets differing by exactly one point of an arbitrary index $i_{0}$; i.e., $\mathbf{x}_{i}=\mathbf{x}_{i}^{\prime}$ for all $i\neq i_{0}$ and $\mathbf{x}_{i_{0}}\neq\mathbf{x}_{i_{0}}^{\prime}$
(12) $Then,$ $\quad\left|\Phi(S)-\Phi(S^{\prime})\right|$
$=\left|\sup(\mathbb{E}_{\mathbf{x}\sim P}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)\right]-\frac{1}{n}\sum_{i=1}^{n}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x}_{i},\mathbf{c})\right)\right])-\right.$
$\quad\left.\sup(\mathbb{E}_{\mathbf{x}\sim P}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x},\mathbf{c})\right)\right]+\frac{1}{n}\sum_{i=1}^{n}\left[\mathcal{L}\left(f_{\omega}(\mathbf{x}^{\prime}_{i},\mathbf{c})\right)\right])\right|$
$\leq\frac{1}{n}\sup_{f_{\omega}\in\mathcal{F}}\left(\left|\mathcal{L}\left(f_{\omega}(\mathbf{x}_{i_{0}},\mathbf{c})\right)-\mathcal{L}\left(f_{\omega}(\mathbf{x}^{\prime}_{i_{0}},\mathbf{c})\right)\right|\right)$
$\leq\frac{u-v}{n}$
In this way, $\Phi(S^{\prime})-\Phi(S)\leq\frac{u-v}{n}$. We could obtain the similar bound $\Phi(S)-\Phi(S^{\prime})\leq\frac{u-v}{n}$. Therefore, for any $\delta>0$, with Lemma 2, at least the probability $1-\delta$:
(13) $\Phi(S)\leq 2\mathcal{R}_{n}(\mathcal{F}_{\omega})+(u-v)\sqrt{\frac{ln(1/\delta)}{2n}}.$
Furthermore, with Lemma 3, we have:
(14) $\Phi(S)\leq 2\sqrt{\frac{2\text{ln}\Pi_{\mathcal{F}}(n)}{n}}+(u-v)\sqrt{\frac{\text{ln}(1/\delta)}{2n}}.$
Based on above proof, we obtain that for any $\delta>0$ and all $f_{\omega}\in\mathcal{F}_{\omega}$, with probability at least $1-\delta$:
(15) $\mathbb{E}_{\mathbf{x}\sim\mathcal{P}}[\ell(f_{\omega}(\mathbf{x},\mathbf{c}))]-\mathbb{E}_{\mathbf{x}\sim\mathcal{P}^{n}}[\ell(f_{\omega}(\mathbf{x},\mathbf{c}))]$
$\leq 2\sqrt{\frac{2\text{ln}\Pi_{\mathcal{F}}(n)}{n}}+(u-v)\sqrt{\frac{\text{ln}(1/\delta)}{2n}}.$
A.11. Limitations & Future Work
In this paper, we propose a novel intent learning method named ELCRec based on the end-to-end learnable clustering framework. It can better mine users’ underlying intents via unifying representation learning and clustering optimization. Besides, the end-to-end learnable clustering module optimizes the clustering distribution via mini-batch data, thus improving the scalability and convenience of deployment. Moreover, we demonstrate the superiority, effectiveness, efficiency, sensitivity, convergence, and visualization of ELCRec on four benchmarks. ELCRec is also successfully applied in the real-world industrial application.
Although achieving promising results, we admit the proposed ELCRec algorithm has several limitations and drawbacks. We summarize them as follows.
  • Pre-defined Cluster Number. The cluster number in the ELCRec method is a pre-defined hyper-parameter. In the real-world data, it is hard to determine the cluster number, especially under the unsupervised conditions. In this paper, for the open benchmarks, we search the cluster number in 32, 64, 128, 256, 512. Besides, for the real-world application, the cluster number is set to 20 based on the number of user groups. However, either the search method or the expert knowledge can not determine the cluster number well at once. The cluster number may change dynamically during model training, and the proposed method may fail to achieve promising performance.
  • Limited Recommendation Domains. In this paper, we adopt four recommendation datasets, including Sports, Beauty, Toys, and Yelp. But, these four datasets are all business recommendation datasets. Besides, in the industrial application aspect, the used data is also a business recommendation dataset. Therefore, the recommendation domain is limited and this paper merely demonstrates the superiority of ELCRec in one domain.
  • No Strategy for Solving Rapid Preference Change. In real-world applications, the intents of users usually change rapidly. Although the proposed ELCRec can capture the user’s intent effectively in the training data, the user’s intent may change dynamically in the future. However, this work does not provide any strategies to alleviate this issue, which may lead to unpromising performance in real-world applications.
To solve the above problems and further improve our method, we summarize several future works and the potential technical solutions as follows.
  • Density-based Clustering. As mentioned above, the cluster number is a pre-defined value in this paper, limiting the recommendation performance and flexibility of the method. To solve this issue in the future, firstly, we can determine the cluster number based on some cluster number estimation methods. They can help to determine the cluster number by performing multiple clustering runs and selecting the best cluster number based on the unsupervised criterion. The mainstream cluster number estimation methods (cluster_number_review, ) include the thumb rule, ELBOW (ELBLOW, ), $t$-SNE (t-SNE, ), etc. The thumb rule simply assigns the cluster number $k$ with $\sqrt{n/2}$, where $n$ is the number of samples. This manual setting is empirical and can not be applicable to all datasets. Besides, the ELBOW is a visual method. Concretely, they start the cluster number $k=2$ and keep increasing $k$ in each step by 1, calculating the WSS (within-cluster sum of squares) during training. They choose the value of $k$ when the WSS drops dramatically, and after that, it reaches a plateau. However, it will bring large computational costs since the deep neural network needs to be trained with repeated times. Another visual method termed $t$-SNE visualizes the high-dimension data into 2D sample points and helps researchers to determine the cluster number. The effectiveness of $t$-SNE heavily relies on the experience of researchers. Therefore, secondly, we can determine the cluster number based on the data density (DPCA, ; DeepDPM, ). Concretely, the areas with high data density are identified as the cluster centers, while the areas with low data density are identified as the decision boundaries between cluster centers. Through this design, the cluster number will be changeable during the training process. It will be determined based on the embeddings itself, better revealing the users’ behavior and may achieve better recommendation performance.
  • More Recommendation Domains. As mentioned above, we admit applied domains of our proposed method are limited. We aim to test our proposed ELCRec on other domains like news recommendation and movie recommendation in the future. For the news recommendation, we can conduct experiments on MIND dataset. This dataset is a large-scale dataset for news recommendation research. It was collected from anonymized behavior logs of the Microsoft News website. MIND contains about 160k English news articles and more than 15 million impression logs generated by 1 million users. For the movie recommendation, we conduct experiments on the MovieLens 1M dataset (ML-1M), Netflix, Douban, Twitch-100k, Twith-full datasets, etc.
  • More Recommendation Domains. As mentioned above, in the real-world application, the intents of the users may change rapidly. However, there is no strategy in the proposed ELCRec method to solve this problem. To this end, in the future, we can propose new cluster center update strategies, such as momentum updates, to control the change rate of the users’ intents. Namely, $\mathbf{C}_{t}=m\cdot\mathbf{C}_{t}+(1-m)\cdot\mathbf{C}_{t-1}$. Here, $\mathbf{C}_{t}$ denote the cluster center embeddings at $t$ and $m$ denotes the momentum. Then, the cluster centers (intents of users) will be changed rapidly when $m$ is large, and the cluster centers (intents of users) will be changed slowly when $m$ is small. This strategy will control the change rate of the users’ intent embeddings, therefore alleviating the above problem.